OpenStudy (anonymous):

Who know the midpoint of D(–5, 2) and E(–13, 0)

OpenStudy (kobeni-chan):

Find the average of each number. -5 + -13 = -18 -18/2 = -9 2 - 0 = 2 2/2 = 1

OpenStudy (kobeni-chan):

(-9,1)

OpenStudy (texaschic101):

add your x's together and divide by 2...add your y's together and divide by 2 (-5,2)...x = -5 and y = 2 (-13,0)...x = -13 and y = 0 add x's and divide by 2.. (-5 + (-13) / 2 = (-5 - 13) / 2 = -18/2 = - 9...thats your x add y's and divide by 2 (0 + 2) / 2 = 2/2 = 1...thats your y so you end up with : (9,1)